On January 1, 2011, Matin Inc. (a wholly-owned subsidiary) bought equipment from Musial (parent) Corp.  for $168,000...

  1. On January 1, 2011, Matin Inc. (a wholly-owned subsidiary) bought equipment from Musial (parent) Corp.  for $168,000 in cash. The equipment originally cost $140,000 but had a book value of only $98,000 when transferred. On that date, the equipment had a five-year remaining life. Depreciation expense was calculated using the straight-line method.


Musial earned $308,000 in net income in 2011 (including investment income) while Matin reported $126,000. Assume there is no amortization related to the original investment. Musial Corp bought inventory constantly from Matin Inc. Three years intra-entity inventory transfer figures are shown in the following table. Assume that gross profit percentage is 20%.

2010

2011

2012

Purchase by Matin

80,000

120,000

150,000

Ending inventory on Matin’s book

12,000

40,000

30,000


Prepare a schedule of consolidated net income and the share to controlling and non-controlling interests for 2011, assuming that Musial owned only 90% of Matin.
Answer #1

NET INCOME CALCULATIONS Martin Inc. gain on sale = 168,000-98,000 = 70,000 Musical corp. depreciation = 168,000/5 = 33,600 CO